OpenStudy (anonymous):

The integral of 2x/sin^2(x^2) dx

OpenStudy (anonymous):

Can you pick a u substitution that might make that a bit easier?

OpenStudy (amistre64):

since the top part is the derivative of the "x" on the bottom, this should just be a log answer. Du(ln(u)) = Du/(u) [S] 2x/[sin(x^2)]^2 dx u = sin(x^2) du = 2x cos(x^2) .... i see the delimma lol

OpenStudy (amistre64):

convert sin^2 to its identity: sin^2 = 1+cos(2x)/2 maybe

OpenStudy (amistre64):

decompose the fraction?